OpenStudy (anonymous):

Given the equation of the parabola The focus of the parabola is: (0, 9) (-9, 0) (0, -9)

OpenStudy (anonymous):

@juliet8

OpenStudy (anonymous):

\[ \large \text{Parabolas}\\ \begin{array}{r|c|c} &\text{Standard}&\text{Vertex}\\ \hline \text{Forms}&y=ax^2+bx+c &y=a(x-h)^2+k\\ \text{Discriminant}& b^2-4ac & -4ak \\ \text{Roots}&x=\frac{-b\pm \sqrt{b^2-4ac}}{2a} & x=h\pm\sqrt{-\frac ka}\\ \text{Vertex}&\left(-\frac b{2a}, c-\frac{b^2}{4a}\right) &\left(h,k\right)\\ \text{Focus}&\left(-\frac b{2a}, c-\frac{b^2-1}{4a}\right) &\left(h,k+\frac{1}{4a}\right)\\ \text{Directrix}&y= c-\frac{b^2+1}{4a} & y=k-\frac{1}{4a}\\ \end{array} \]

OpenStudy (anonymous):

so its the second one?

OpenStudy (anonymous):

Another equation for a parabola is \(4py=x^2\).

OpenStudy (anonymous):

In this case the focus is at \((0,p)\).

OpenStudy (anonymous):

It turns out that \(4p = 1/a\).

OpenStudy (anonymous):

in our case \(4p = -36\).

OpenStudy (anonymous):

Find \(p\).

OpenStudy (anonymous):

(0,-9)?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

so the directrix of the parabola is -9? @wio

OpenStudy (anonymous):

Directrix is \(y=-p\).

OpenStudy (anonymous):

so \(y=-(-9) = 9\)

OpenStudy (anonymous):

thnx:)